# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws [16]. In addition to increasing the depth of neural network models, widening neural networks by using the sparsely-activated mixture-of-experts (MoE) architecture is also proved effective. MoE widens the feed-forward network (FFN) layer (one expert) by having multiple parallel FFNs (experts). During forward propagation, only a subset of these experts is activated. Thus, compared to dense models, MoE models achieve better end-task performance and generalize better to new tasks without increasing computation costs. Notable examples of MoE models include Switch Transformer [9], Mixtral-MoE [14], and Uni-MoE [22].
@@
Despite significant progress in developing wider and deeper MoE LLMs, the increased memory consumption due to larger model sizes (i.e., increased number of experts) poses a substantial challenge to the deployment of these models in real-world settings. For example, storing and loading Mixtral-8x7B, which has 8 experts in each of its 32 layers, requires approximately $88$ GB. The MoE layers constitute the majority of the parameters. Adding or removing even one expert in each layer can significantly impact overall memory cost and model performance. For example,  [26] shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark [12]. In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
@@
There have been several studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. [3] prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. [5] find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. [21] merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. [11] explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. [26] enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
@@
Figure 1: Removing several experts from the original MoE layer would not cause model collapse but improve efficiency. Prior works often utilize expert access information to filter out unimportant experts. In our work, we first group different experts with similar knowledge in the feature space, then merge them along with the routers to prune the MoE layer. This post-processing approach allows us to diversify the features of each MoE layer, thereby preserving the knowledge of the original large models as much as possible while reducing computation and storage consumption.
@@
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 1, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A). The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks.
@@ 2 Related Work
Sparse MoEs. The Mixture-of-Experts (MoE) structure is firstly applied in classical machine learning models by Jacobs et al. [13] and Jordan and Jacobs [15], then widely used in various deep learning models [39, 29, 40]. Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models [32, 19, 43]. It adapts the original large feed-forward network (FFN) in each transformer block into multiple smaller FFNs, forming an expert layer with a router that computes the weighted output of each MoE layer. Sparse MoEs were first proposed by Fedus et al. [9]. In this approach, only a few experts are activated in each layer, accelerating training and inference while significantly increasing the number of parameters for greater model capacity. Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer [9], Mixtral-8B [14], and Uni-MoE [22]. Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models [41, 20].
Model Pruning. Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency [25, 36, 24]. Pruning techniques can be categorized into unstructured pruning [23, 34, 28], which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning [18, 8, 33], which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs [4]. Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs [27], RNNs [42], and LSTMs [7] to modern large models such as Llama [38] and Stable-Diffusion [2].
@@ 3.1 Task Definition
Previous works [26, 11, 21] have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
@@ 3.2 Expert Similarity in MoE layer
We employ the Centered Kernel Alignment (CKA) [17] as the criteria to evaluate the similarity between experts in each MoE layer. Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
@@
Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 [31]. We respectively use the linear kernel and RBF kernel in similarity computation of eq. 4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
As illustrated in Fig. 2, the darker the cell, the more similar the two related experts. The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers. There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts. For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance.
@@ 3.3 Discover and Merge Similar Experts on Feature Space
Figure 3: Overview of our method. We first leverage different metrics to obtain the expert representation, then compute the pairwise distance to get the expert similarity matrix. Based on the expert similarity matrix, we can group the experts with shared knowledge in the same cluster, which can be merged on the weight space for MoE pruning.
@@
Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the distance matrix $D$.
@@
Merging Similar Experts with Routers. To diversity the experts and preserve the different knowledge learned by different models clustered in the same group [37, 35], we merge the clusted experts with their routers on the weight space as follows,
@@ 3.4 Practical Implementation
For data-centric strategy, we follow previous work [26] to use the pre-trained dataset C4, which can be a good surrogate to simulate the performance of the task-agnostic pruning. To improve the generalization ability at the discovery stage, we randomly mix the token embedding for data augmentation and then use the augmented data to calculate the expert output as the expert representation.
For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation. The first way is to leverage the vectorized weights directly. The other way is to leverage the local linearity of neural networks to compute the surrogate weight matrix.
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$. Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance in pruning models (see section A.2).
@@
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups. Among these three strategies, we empirically find that the uniform souping strategy offers more stable performance and greater efficiency in model pruning (see section A.2 ). While the learning strategy can yield slightly better performance, it is more time-consuming due to the need for tuning the parameter $\alpha$.
@@ 4 Evaluation
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Router-guided [21] 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
Count-guided [11] 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
Enumerate [26] 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
@@ 4.1 Experiment Setup
Pruning Methods. We take three advanced MoE pruning methods [11, 21, 26] as our baseline for comparison. Among them, router-guided merging [21] is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming [26], also the count-guided strategy and loss-based pruning method [26], namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
For our method, we respectively report the best results of our data-centric and model-centric methods in pruning models. Following the previous setting [26], we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data. More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in section A.2.
@@
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Router-guided [21] 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
Count-guided [11] 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
Enumerate [26] 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
@@
Evaluation Datasets. The open-sourced Language Model Evaluation Harness library [10] is used to evaluate the performance. We select four zero-shot tasks, including MMLU [12], BoolQ [6], OpenBookQA [30], and RTE [1]. Among these tasks, MMLU consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have achieved outstanding performance across various tasks by learning a large number of model parameters on large amounts of data, as shown by the scaling laws (Kaplan et al., 2020). In addition to increasing the depth of neural network models, widening neural networks by using the sparsely-activated mixture-of-experts (MoE) architecture is also proved effective. MoE widens the feed-forward network (FFN) layer (one expert) by having multiple parallel FFNs (experts). During forward propagation, only a subset of these experts is activated. Thus, compared to dense models, MoE models achieve better end-task performance and generalize better to new tasks without increasing computation costs. Notable examples of MoE models include Switch Transformer (Fedus et al., 2022), Mixtral-MoE (Jiang et al., 2024), and Uni-MoE (Li et al., 2024b).
@@
Despite significant progress in developing wider and deeper MoE LLMs, the increased memory consumption due to larger model sizes (i.e., increased number of experts) poses a substantial challenge to the deployment of these models in real-world settings. For example, storing and loading Mixtral-8x7B, which has 8 experts in each of its 32 layers, requires approximately $88$ GB. The MoE layers constitute the majority of the parameters. Adding or removing even one expert in each layer can significantly impact overall memory cost and model performance. For example,  (Lu et al., 2024) shows that randomly dropping $2$ experts in each MoE layer reduces the memory cost by $21$ GB, and decreases model performance by $7\%$ on the MMLU benchmark (Hendrycks et al., 2020). In this study, we strive to seek the best trade-off between memory efficiency and task performance by identifying an optimal set of experts in each MoE layer to prune.
@@
There have been several studies on pruning MoE models. One line of work utilizes task-specific information to prune irrelevant experts. For example, Chen et al. (2022) prune the less frequently visited experts based on experiments on a range of tasks. Chowdhury et al. (2024) find that less important experts usually exhibit smaller changes in routing weights during the fine-tuning stage. Li et al. (2024a) merge experts that are frequently visited by tokens of a fine-tuned dataset for pruning. Compared with task-specific pruning methods, task-agnostic pruning is more challenging due to the lack of explicit criteria to evaluate the redundancy of experts on unknown tasks. He et al. (2024) explore pruning experts with less visited frequency in a task-agnostic calibration dataset but report a significant performance drop. In comparison, Lu et al. (2024) enumerate all the combinations of experts and prune some to achieve a minimum loss of reconstruction, which significantly improves performance. We illustrate the difference among these works in fig. 1. Although pruning MoE models in task-agnostic settings is of great practical value, this area has not been fully explored.
@@
Figure 1: Removing several experts from the original MoE layer would not cause model collapse but improve efficiency. Experts with similar colors share the similar knowledge with each other. Prior works often utilize expert access information to filter out unimportant experts. In our work, we first group different experts with similar knowledge in the feature space, then merge them along with the routers to prune the MoE layer. This post-processing approach allows us to diversify the features of each MoE layer, thereby preserving the knowledge of the original large models as much as possible while reducing computation and storage consumption.
@@
In this work, we explore how to prune MoE models in a task-agnostic fashion. Our study is motivated by the finding that, given the same input, many experts respond similarly, indicating that these experts likely encode similar knowledge, and thus are somewhat redundant. We propose a method to improve model parameter efficiency by pruning redundant experts in two stages. As shown in fig. 3, we first identify and group similar experts in the feature space. Then, for each group, we merge experts in the weight space to diversify the knowledge in different MoE layers. We validate the effectiveness of our method by pruning experts for three state-of-the-art MoE architectures, including Mixtral (Mixtral-8x7B and Mixtral-8x22B), Deepseek-MoE (Deepseek-MoE-16B) and Qwen (Qwen2-57B-14A). The evaluation shows that our method outperforms other model pruning methods on a range of natural language tasks. Our contribution can be summarized as follows,
  1. 1.
    We empirically validate that some experts within each well-trained MoE layer encode similar knowledge, making them somewhat redundant.
  2. 2.
    We propose a two-stage, task-agnostic method for grouping and merging redundant experts, which is further divided into data-centric and model-centric implementation strategies.
  3. 3.
    We demonstrate the effectiveness of our method by pruning experts from a series of state-of-the-art MoE models, including Mixtral-MoE, DeepSeek-MoE, and Qwen-MoE. The results from a greedy search for MoE pruning further validate the success of our approach.
@@ 2 Related Work
Sparse MoEs. The Mixture-of-Experts (MoE) structure is firstly applied in classical machine learning models by Jacobs et al. (1991) and Jordan and Jacobs (1994), then widely used in various deep learning models (Yuksel et al., 2012; Masoudnia and Ebrahimpour, 2014; Zhang et al., 2023). Recently, some works employ the MoE to scale the capacity of transformer-based models, especially the large language models (Shazeer et al., 2017; Lepikhin et al., 2020; Zoph et al., 2022). It adapts the original large feed-forward network (FFN) in each transformer block into multiple smaller FFNs, forming an expert layer with a router that computes the weighted output of each MoE layer. Sparse MoEs were first proposed by Fedus et al. (2022). In this approach, only a few experts are activated in each layer, accelerating training and inference while significantly increasing the number of parameters for greater model capacity. Many sparse MoEs have been developed and open-sourced within the AI community, such as Switch Transformer (Fedus et al., 2022), Mixtral-8B (Jiang et al., 2024), and Uni-MoE (Li et al., 2024b). Recent studies also indicate that neural networks with the MoE structure exhibit better generalization ability compared to dense models (Zhu et al., 2022; Li et al., 2022).
Model Pruning. Model pruning involves removing unimportant parameters from a well-trained neural network to balance task performance and computational efficiency (Liu et al., 2018; Wang et al., 2020; Liu et al., 2021). Pruning techniques can be categorized into unstructured pruning (Liao et al., 2023; Shi et al., 2024; Mason-Williams and Dahlqvist, 2024), which introduces sparsity in the weight matrix by setting some parameters to zero, and structured pruning (Lemaire et al., 2019; Fang et al., 2023; Shen et al., 2022), which removes entire neurons, layers, or blocks, reducing redundancy and being more suitable for acceleration on GPUs (Choi and Yang, 2021). Many efforts have been made to leverage model pruning techniques to reduce the memory consumption of neural networks, spanning a range of models from conventional architectures like CNNs (Luo et al., 2018), RNNs (Zhu and Gupta, 2017), and LSTMs (Ding et al., 2020) to modern large models such as Llama (Xia et al., 2023) and Stable-Diffusion (Castells et al., 2024).
@@ 3.1 Task Definition
Previous works (Lu et al., 2024; He et al., 2024; Li et al., 2024a) have identified that removing some experts from different MoE layers does not incur model collapse but improves efficiency in inference and fine-tuning on downstream tasks. The remaining question is which expert(s) should be pruned in the MoE layer.
@@ 3.2 Expert Similarity in MoE layer
We employ the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) as the criteria to evaluate the similarity between experts in each MoE layer. Specifically, given a batch of inputs $\{x_{1},x_{2},...,x_{s}\}$, the similarity $\rho_{ij}$ between two experts $f_{i}$ and $f_{j}$ is computed as follows,
@@
Here, we compute the expert similarity of Mixtral-8Bx7B using randomly selected $32$ samples from the pre-trained dataset C4 (Raffel et al., 2020). We respectively use the linear kernel and RBF kernel in similarity computation of eq. 4, where the use of linear kernel actually corresponds to the dot product similarity between the expert outputs.
As illustrated in Fig. 2, the darker the cell, the more similar the two related experts. The two kernel-based distance measures demonstrate similar changes in expert similarity across different layers. There are varying moderate ratios of similar paired experts in different MoE layers, suggesting the plausibility of pruning the MoE layer with many experts down to fewer experts. For example, in the first layer, both the linear kernel and the RBF kernel indicate that the $2$-nd to $5$-th experts exhibit great similarity (greater than 0.7) with each other, suggesting that one of them can be pruned without significantly degrading model performance. Full evaluation results can be found in appendix B.
@@ 3.3 Discover and Merge Similar Experts on Feature Space
Figure 3: We first leverage model or data-centric strategies to obtain the expert representation, then compute the pairwise distance to get the disparity matrix. Based on the expert similarity matrix, we can group similar experts with shared knowledge in the same cluster, which can be merged on the weight space for pruning.
@@
Then, we initialize $\mathcal{G}$ as a complete graph, namely the expert graph, where $\mathcal{V}$ is set as all expert indices, and $\mathcal{E}$ is set as the pair-wise expert distance matrix $D$.
@@
Merging Similar Experts with Routers. To diversity the experts and preserve the different knowledge learned by different models clustered in the same group (Wortsman et al., 2022; Wang et al., 2022), we merge the clusted experts with their routers on the weight space as follows,
@@ 3.4 Practical Implementation
For the data-centric strategy, since the full pre-training datasets for different models are large and inaccessible, we use the C4 dataset, a commonly used smaller pre-training subset that serves as an effective surrogate for capturing task-agnostic knowledge of experts Lu et al. (2024). Specifically, by disabling the router function, for the same input, we use the output of each expert as its expert representation $\mathcal{M}(\cdot)$. To enhance generalization and mitigate overfitting to the selected samples during model pruning, we apply data augmentation by randomly mixing token embeddings during the representation computation at the discovery stage.
For model-centric strategy, we propose two ways to prune models with only expert weights, which encode the dataset information during the training process and can be a good agency for expert representation. One is to leverage the vectorized weights directly ($\clubsuit$). The other is to leverage the local linearity of neural networks (Zhang and Wu, 2020) to compute the surrogate weight matrix ($\spadesuit$).
Taking the FNN $f(\cdot;\theta)$ in Mixtral-8x7B as an example, it consists of three linear layers, i.e., $\theta=\{\theta_{1},\theta_{2},\theta_{3}\}$. For input $x$, the output is computed as $f(x;\theta)=\theta_{2}(\sigma(\theta_{1}x)\cdot\theta_{3}x)$. Then, the vectorized weight ($\clubsuit$) is $\text{concat}\{\theta_{1},\theta_{2},\theta_{3}\}$, while the surrogate weights can be obtained by $\theta_{2}(\theta_{1}\cdot\theta_{3})$ ($\spadesuit$). Compared to vectorized weights, surrogate weights are more flexible with model size, showing more stable performance (see appendix D).
@@
where the ground-truth is the output of original MoE layer $y=F(x;\Theta,W,K)$, and we jointly optimize $\lambda$ and $\alpha$ for different merging groups. Among these three strategies, both the first and third require the presence of data, while the second is compatible with both data-centric and model-centric approaches. We empirically find that the uniform souping strategy offers more stable performance and greater efficiency in task-agnostic model pruning (see appendix D ). While the learning strategy can yield slightly better performance, it is more time-consuming due to the need for tuning the parameter $\alpha$.
@@ 4 Evaluation
Mixtral-8x7B (Jiang et al., 2024) 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Router-guided (Li et al., 2024a) 51.8/24.8 60.5/26.5 46.9/24.7 60.5/25.0 82.6/39.9 32.0/11.6 70.4/50.9 57.8/29.1
Count-guided (He et al., 2024) 49.2/36.9 59.7/45.6 45.0/35.1 58.2/43.4 77.2/76.6 33.0/26.4 56.6/55.9 54.1/45.7
Enumerate (Lu et al., 2024) 52.4/43.5 66.4/52.7 49.0/40.4 63.7/43.5 84.0/80.8 32.6/28.8 71.1/66.4 59.9/50.8
@@ 4.1 Experiment Setup
Pruning Methods. We take three advanced MoE pruning methods (He et al., 2024; Li et al., 2024a; Lu et al., 2024) as our baseline for comparison. Among them, router-guided merging (Li et al., 2024a) is initially designed for task-specific MoE pruning, where we set the target dataset as the samples from the pre-training dataset. For task-agnostic methods, we select the frequency-based pruning method Expert Trimming (Lu et al., 2024), also the count-guided strategy and loss-based pruning method (Lu et al., 2024), namely the Enumerate in our paper. Under the task-agnostic pruning setting, we disable all the fine-tuning stage of all methods for fair comparison.
For our method, we respectively report the best results of our data-centric and model-centric methods in pruning models. Following the previous setting Lu et al. (2024), we use $128$ samples in C4 for computation in data-centric pruning methods, while the model-centric method doesn’t rely on the data. More detailed results on different implementations of discovery and merging steps (e.g., vectorized and surrogate weight strategy at the discovery step, max or learning strategy at the merging step) are provided in appendix D.
@@
Mixtral-8x22B (Jiang et al., 2024) 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Router-guided (Li et al., 2024a) 27.3/22.7 25.4/25.8 24.4/24.0 27.9/23.4 62.8/62.7 12.8/13.0 54.2/49.5 33.5/31.6
Count-guided (He et al., 2024) 58.0/45.7 74.9/57.7 54.1/42.0 70.2/45.7 81.5/74.4 35.2/27.0 69.3/57.4 63.3/50.0
Enumerate (Lu et al., 2024) 60.4/53.9 78.0/67.2 59.5/52.3 73.0/64.2 87.4/80.5 35.0/31.1 70.1/67.9 66.2/59.6
@@
Evaluation Datasets. The open-sourced Language Model Evaluation Harness library (Gao et al., 2021) is used to evaluate the performance. We select four tasks, including MMLU (Hendrycks et al., 2020), BoolQ (Clark et al., 2019), OpenBookQA (Mihaylov et al., 2018), and RTE (Bentivogli et al., 2009). Among these tasks, MMLU is the most challenging one, which consists of $57$ subtasks, where we present four groups, namely the humanities, social science, stem and other.